Suppose you have a six sided die. One face is printed with the number 1. Two faces are printed with the number 2. Three facesPlease answer all parts to this 4 part question
Answer #1

ANSWER:

1)

Given

P(C1) = 1/6
P(C2) = 2/6
P(C3) = 3/6

Also, we are given that:
P(heads | C1) = 1/5
P(heads | C2) = 1/3
P(heads | C3) = 1/2

Using law of total probability, we get here:
P(heads) = P(heads | C1)P(C1) + P(heads | C2)P(C2) + P(heads | C3)P(C3)

P(heads) = (1/5)*(1/6) + (1/3)*(2/6) + (1/2)*(3/6) = 0.3944

Therefore 0.3944 is the required probability here.

2)

Given the toss is heads, probability that the die landed on a 2 is computed as:

P(coin_2|heads) = P(2 on die and heads) / P(heads)

\small =\frac{(2/6)*(1/3)}{0.3944}

= 0.2817

Therefore 0.2817 is the required probability here.

3)

P(winning first on 3rd attempt) = (1 - 0.2817)3 * 0.2817

= 0.1044

Therefore 0.1044 is the required probability here.

4)

The expected winnings here is computed as:

=  1(1/5)*(1/6) + 2(1/3)*(2/6) + 3(1/2)*(3/6)

= 0.0333 + 0.2222 + 0.75

= 1.005

Therefore expected winnings = 1.005

If you have any doubts please comment and please don't dislike.

PLEASE GIVE ME A LIKE. ITS VERY IMPORTANT FOR ME.